OpenStudy (toxicsugar22):

how do u solve sqrt(x+4)=sqrt(x)+1

OpenStudy (toxicsugar22):

and show me the steps

myininaya (myininaya):

\[\sqrt{x+4}=\sqrt{x}+1\] square both sides \[(\sqrt{x+4})^2=(\sqrt{x}+1)^2\] \[x+4=\sqrt{x}\sqrt{x}+2\sqrt{x}+1\] \[x+4=x+2\sqrt{x}+1\] subtract x on both sides also subtract 1 on both sides \[x-x+4-1=2\sqrt{x}\] \[3=2\sqrt{x}\] divide both sides by 2 \[\frac{3}{2}=\sqrt{x}\] now square both sides \[(\frac{3}{2})^2=(\sqrt{x})^2\] \[\frac{9}{4}=x\]

OpenStudy (anonymous):

very nice

OpenStudy (anonymous):

like this one?

myininaya (myininaya):

aww your so sweet :) its the same as the last one

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

this is a car

myininaya (myininaya):

its still the red one lol

OpenStudy (anonymous):

the last one was a bike

OpenStudy (toxicsugar22):

can u stop typing please

myininaya (myininaya):

i didn't see a bike

OpenStudy (toxicsugar22):

can help me with one more

OpenStudy (toxicsugar22):

solve the equation 3^(x+6)=4

OpenStudy (anonymous):

\[3^{x+6}=4\] \[(x+6)\ln(3)=\ln(4)\] \[xln(3)+6\ln(3)=\ln(4)\] \[x\ln(3)=\ln(4)-\ln(6)\] \[x=\frac{\ln(4)-\ln(6)}{\ln(3)}\]

OpenStudy (anonymous):

i made a mistake

myininaya (myininaya):

lol you did 6ln3 does not equal ln6

OpenStudy (anonymous):

should be \[x\ln(3)=\ln(4)-6\ln(3)\]

OpenStudy (anonymous):

so \[x=\frac{\ln(4)-6\ln(3)}{\ln(3)}\]

myininaya (myininaya):

satellite im gonna give you medals for everything we are gonna get you built up i'm gonna go om amistre medal strike

OpenStudy (anonymous):

you can combine the logs in the numerator if you like but i wouldn't bother

OpenStudy (anonymous):

i will never catch up but thanks anyway.

myininaya (myininaya):

you can write it has two separate fractions and cancel the ln3 from the last fraction

OpenStudy (anonymous):

now it is beer o'clock. later

myininaya (myininaya):

\[x=\frac{\ln4}{\ln3}-6\]